In an LC oscillating circuit, the maximum charge on capacitor is Q 0 . The current in the circuit at an instant when energy stored in inductor is equal to that in capacitor, is

In an LC oscillating circuit, the maximum charge on capacitor is Q0. The current in the circuit at an instant when energy stored in inductor is equal to that in capacitor, is
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
Total energy
U= 1/2(Q^2​/C)
Let at any instant the charge on capacitor beQ’
∴Electrical energy = Magnetic energy= ((Q’)^2​)/2C
So, new total energy=((Q’)^2​)/2C +((Q’)^2​)/2C =Q′^2​/C
According to question
1/2(Q^2​/C)=Q′^2​/C
Q′= Q/​√2
Magnetic energy= ((Q’)^2​)/2C= Q^2/4C
(1/2)L(I^2) =Q^2/4C

I = Q / √(2LC)


Thanks